If two forces of 14 Newtons act on an object, one up and to the right at 45 degrees, and the other up and to the left in 45 degrees, then the net sideways force is?
1
Expert's answer
2019-09-26T14:48:39-0400



The net force is directed upward


"F_{\rm net}=F_1\cos 45^{\circ}+F_2\cos 45^{\circ}""=14\:\rm N\cos 45^{\circ}+14\:\rm N\cos 45^{\circ}=19.8\:\ N"

The net sideways force is zero.